# Local Spherical Collapsing Box in Athena++: Numerical Implementation and Benchmark Tests

### v1

@@ 4.2.1 Horizontal soundwaves with uniform collapse
We first consider a purely horizontally propagated traveling wave with a uniform collapse profile. This can also provide a comprehensive view of the wave behavior in the local box. We set up the following perturbation
@@
where $A_{0}$ is some (real) constant, $c_{\mathrm{s},\mathrm{H0}}=c_{\mathrm{s0}}/R_{0}$ is the initial value of the effective horizontal sound speed as defined in Section 2. This perturbation corresponds to wave amplitudes of $\left(X,\Pi\right)=\left(\frac{iA_{0}}{2c_{\mathrm{s}}},\frac{A_{0}k_{x}}{2R_{0}}\right)$.
For the numerical test of the local model, we choose a collapse profile with $R_{0}=10.0L_{0}$, $U_{0}=1.0c_{\mathrm{s}0}$, $L_{z0}=1.0L_{0}$, $U_{R0}=0$, with perturbation amplitude of $A_{0}=10^{-6}$, $k_{x}=16\pi$. The simulation box size is set as $L_{x,\mathrm{sim}}=0.5L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 256 cells per unit length.
The relative amplitude of gas density perturbation $\delta\rho/\rho$ is measured with respect to the background gas density $\rho$ at each time point, which is obtained by averaging the gas density over the domain. The space-time plot of the vertically averaged value of $\delta\rho/\rho$ presented in the upper panel of Figure 5 shows the evolution of the wave pattern. The wave propagates rightward with a time-varying effective sound speed $c_{\mathrm{s},\mathrm{H}}=c_{\mathrm{s}0}/\mathcal{R}$. As the collapse proceeds, the wave propagation is accelerated as $\mathcal{R}$ decreases, matching the overlayed curves of theoretical phase shift profiles over time. This time-varying effective sound speed is not physical in the global picture, but is instead due to our choice of a time-dependent reference frame in the local model.
Our choice of the collapse profile results in the wave operating in the WKB regime. It is thus appropriate to compare the evolution of the wave amplitude to that expected from the WKB approximation solution. The lower panel of Figure 5 shows this comparison of the relative amplitude of density perturbation, $A_{\rho}=\max\left(\delta\rho/\rho\right)$. The numerical result of the $A_{\rho}$ evolution agrees well with the WKB approximation solution, $A_{\rho}=A_{\rho,0}\mathcal{R}^{-1/2}$.
Similarly, for a purely vertical sound wave, the corresponding length scale becomes $L_{z}$ instead of $\mathcal{R}$. We have also numerically tested the vertical sound wave and found the effective vertical sound speed evolves following $c_{s,H}=c_{s0}/{L_{z}}$, and the relative amplitude of $\delta\rho$ follows $A_{\rho}=A_{\rho,0}L_{z}^{-1/2}$, which aligns with the theoretical expectations.
4.2.2 Horizontal Waves with non-uniform collapse
We now test the horizontal sound waves with a more general power law collapse profile. In order to compare with the exact solution Eq. (90-92), we consider a collapse profile with $\beta=1/2$ and $t_{\mathrm{c}}=1$, i.e., $\mathcal{R}={R}_{0}(1-t)^{1/2}$, and initiate correspondingly the perturbation of velocity and density as
@@ 4.2.2 Horizontal Waves with non-uniform collapse
where according to eq. (93), $s_{0}=s(0)=\left(t_{c}/t_{0}\right)^{1-\beta}$, and $t_{0}$ can again be obtained from eq. (93). $A_{0}$ is some real constant, and this perturbation corresponds to $A=0$, $B=A_{0}/2$ in eq. (90-92). For the numerical test, we set $R_{0}=1.0L_{0}$, $L_{z0}=1.0L_{0}$, $U_{R0}=0$, with perturbation of $A_{0}=10^{-5}$, $k_{x}=2\pi$, and simulation box size $L_{x,\mathrm{sim}}=2.0L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 256 cells per unit length.
Similar to §4.2.1, the relative amplitude of gas density perturbation $\delta\rho/\rho$ is measured at each time point, where $\rho$ is obtained by averaging the gas density over the domain. The space-time plot of vertically averaged relative gas density $\delta\rho/\rho$ displayed in the upper panel of Figure 6 shows that this perturbation leads to a horizontal standing wave. This wave exhibits phase reversal with each half wave period. We further measure the relative amplitude of the density perturbation by performing a Fourier transform to the wave pattern, and identify the amplitude of the dominant frequency component. The time evolution of the relative amplitude $\delta\rho/\rho$ is shown in the lower panel of Figure 6, with the length scale corresponding to freeze-out regime $L_{\mathrm{freezeout}}$ (see Eq. (94)) marked with the vertical dotted line. The result is in good agreement with the analytical solution in both the WKB regime ($\mathcal{R}>L_{\mathrm{freezeout}}$) and the freeze-out regime ($\mathcal{R}<L_{\mathrm{freezeout}}$). The L2 norm of the relative error for the relative amplitude $\mathrm{err}(\delta\rho/\rho)$ as defined in eq. (70) is of the order of $10^{-5}$ when the absolute value of the amplitude solution is much larger than 00. When the solution of the amplitude is close to 00 (i.e. close to the period of phase reversal), the relative error $\mathrm{err}(\delta\rho/\rho)$ can be up to the order of $10^{-2}$, partly due to the division of a small analytical solution value, and partly due to inherent limitations and numerical inaccuracies that can arise near the phase reversal period. The slightly diffused wave pattern during the phase reversal can also be seen in Fig. 6.
Figure 7: Test case for diagonal sound wave. Upper panel: illustration of the diagonal sound wave and the shear flow generated; Lower panel: curve of vortical wave growth as a test parameter. Lower panel: amplitude of the shear flow generated by the wave as a function of R. The numerical result (red solid line), closely matches the analytical solution (black dashed line).
@@ 4.2.3 Diagonal soundwaves
Diagonally propagating sound waves may introduce numerical complexities, as they generate shear flows perpendicular to the direction of wave propagation. This is illustrated in the upper panel of Figure 4, with the wave pattern characterized by the density.
We test diagonal sound waves by generating density and velocity perturbations of
@@
where $A_{0}$ is some (real) constant. This perturbation corresponds to an initial wave amplitude of $\left(X_{0},\Pi_{0}\right)=\left(\frac{iA_{0}}{2c_{\mathrm{s}0}},\frac{A_{0}}{2}\sqrt{k_{x}^{2}R_{0}^{-2}+k_{z}^{2}L_{z0}^{-2}}\right)$. Recall Eq. (80), Eq. (81) and Eq. (95), this pure diagonal sound wave generates a shear flow that can be characterized by
@@
and the dynamics of the system of $(X,\Pi)$ can be obtained from Eqs. (80) and (81). Thus the amplitude $X$ can be obtained by solving the equation
@@
where the wave frequency $\omega$ is given in Eq. (82).
We numerically test the diagonal sound wave by setting $R_{0}=1.0L_{0}$, $U_{0}=1.0c_{\mathrm{s}0}$, $L_{z0}=1.0L_{0}$, $U_{R0}=0$, with perturbation $A_{0}=10^{-5}$, $k_{x}=2\pi$, $k_{z}=4\pi$, and simulation box $L_{x,\mathrm{sim}}=2.0L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 512 cells per unit length.
In the lower panel of Fig. 7, we compare the amplitude of the shear flow generated by the diagonal soundwave in the numerical simulation with the analytical solution. The numerical result is obtained by calculating the shear flow $\delta v_{\mathrm{shear}}$ in each grid, and performing a 2D Fourier Transform of the shear flow map, followed by identifying the amplitude of the dominant frequency component. The analytical solution of the shear flow amplitude $2{c_{\mathrm{s}}\mathcal{R}^{-2}}k_{x}k_{z}\left(1-b^{-2}\right)X$ is obtained by numerically solving for the wave amplitude $X$ at each time step according to Eq. (108), with the explicit Runge-Kutta integration method of order 5(4) (Dormand & Prince, 1980). The numerical result and analytical solution of the shear flow amplitude are in good agreement. We further measure the L2 norm of the relative error for the shear strength $\delta v_{\mathrm{shear}}$ in each grid, and get $\mathrm{err}(\delta v_{\mathrm{shear}})$ as defined in Eq. (70) of the order of $10^{-6}$.
Figure 8: Conservation laws for tests of horizontal shear flow (left) and of white noise in velocity field (right). For each case, the normalized variation of $J\bar{\rho}$ (left), $\Pi_{i}$ (middle) and $H_{k}$ (right) are presented as a function of time. For the horizontal shear flow test (see also §4.1.1), the modified mass conservation law ($J\bar{\rho}$), the conservation law for total momenta ($\Pi_{i}$), and for kinetic helicity $H_{k}$ are all satisfied to machine precision (note that for all of the panels, the y axis ticks are in the unit of $10^{-13}$). For the case with white noise in velocity field, the conservation laws for both mass $J\bar{\rho}$ and total momenta $\Pi_{i}$ are fullfilled to machine precision as well, but not for the kinetic helicity $H_{k}$ (note that for the upper two panels, the y axis ticks are in the unit of $10^{-14}$). For comparison, a simulation with the same setup of white noise, but without collapse (i.e. $U_{0}=U_{R0}=0$) is performed (“WN0”, grey dashed lines, see also the legend), which gives consistent results to WN. We thus anticipate that the deviation from conservation law for $H_{k}$ is likely due to intrinsic characteristics of Athena++, which in certain situations is not strictly vorticity conserving.
@@ 4.3 Conservation Laws
The continuity equation of the local model, Eq. (5), gives the conservation law for the total mass
@@
Momentum equation of the local model, Eq. (6 - 8), gives the conservation law for the total momenta
@@
Another conserved quantity in the local model is the kinetic helicity $H_{k}$ of the relative flow, which is a quantity associated with the vorticity. It is defined as
@@
In order to validate these conservation laws in the local box, we perform a simulation run with uniform background density $\rho_{0}=1$, introducing white noise perturbations in the velocity filed $v_{x},v_{y}$ and $v_{z}$, with a relative amplitude of $10^{-6}$. We choose the collapse profile with $R_{0}=10.0$, $U_{0}=-1.0c_{\mathrm{s}0}$, $L_{z0}=10.0$, $U_{R0}=-0.1$, and set the simulation box size $L_{x,\mathrm{sim}}=L_{y,\mathrm{sim}}=0.5$, $L_{z,\mathrm{sim}}=1.0$, with a resolution of 32 cells per unit length. This test run is labeled as “WN”. For comparison, we also perform a simulation with the same setup of the white noise, but without collapse (i.e. $U_{0}=U_{R0}=0$), which is labeled as “WN0”. In addition, the result for the horizontal shear flow test (see §4.1.1) is also used for examining the conservation laws.
We measure the volume-averaged density $\bar{\rho}$ , total momenta $\Pi$ and kinetic helicity $H_{\rm k}$ in the simulations, and present in Figure 8 the normalized variation of $J\bar{\rho}$ , $\Pi$ and $H_{\rm k}$ as a function of time. Our results demonstrate that the modified mass conservation law ($J\bar{\rho}$) and the conservation law for total momenta ($\Pi_{i}$) are both satisfied to machine precision. The kinetic helicity conservation, although preserved to machine precision for the shear flow test, exhibits significant deviation for the white noise simulations. This discrepancy persists in both WN and WN0 cases, suggesting that the lack of conservation is not a consequence of the integration of the collapsing box. Instead, it is likely due to intrinsic characteristics of the Athena++ framework, which in certain situations is not strictly vorticity conserving, likely because the higher-order Godunov scheme generally introduces a degree of numerical diffusion (see e.g. Seligman & Laughlin, 2017). In the white noise case where vorticity is dominated by fine-scale structures, the increased numerical dissipation in the small scales because the slope limiter is expected to lead to the non-conservation of $H_{\rm k}$. For an adiabatic EOS, the entropy conservation law is also validated, which is further presented in App. C.
@@ 4.4 Convergence Study
We perform a convergence study with the diagonal wave (§4.2.3) test, as it shows the maximum possible error. Basic simulation setups such as wave properties and box sizes are the same as presented before. Fiducially, simulations are performed with Roe’s solver, the VL2 integrator, PPM reconstruction, and isothermal EOS. The CFL number in the convergence study is fixed to $0.2$. However, we have validated the numerical stability of our code for CFL numbers up to 0.4 for VL2 and 0.8 for RK3. We further explore the effects of different Riemann solvers, time integrators, spatial reconstruction methods, and the EOS on numerical performance of the code. We also vary the spatial resolution of our simulation in the unit of cells per unit length. The simulation setups for test runs that we perform for the convergence study are presented in Table 1. Figure 9 shows the time-averaged L2 relative error $\mathrm{err}(\delta v_{\mathrm{shear}})$ as a function of resolution.
@@
Table 1: Simulation setup for the convergence study, with variations compared to the fiducial simulation highlighted in bold. “Reconst.” stands for reconstruction.
Our results demonstrate that our code generally achieves convergence between first- and second-order for the diagonal wave test. The performances for test cases HLLE, and RK3 remain generally similar to the fiducial case, while a lower order of spatial reconstruction (PLM) leads to lower accuracy of lower resolutions. These tests with isothermal EOS achieve second-order convergence at lower resolutions, with a tentative trend towards first-order for high resolutions. This trend is likely attributed to time-dominated errors at higher resolutions, as our source term incorporation is first-order in time. We also explored the higher-order method to integrate source terms with the VL2 integrator, which was not as efficient as our first-order method, showing comparable or slightly higher error, and reduced stability for adiabatic flows. Therefore, we choose to use our first-order method. The source term can also be incorporated into higher-order methods with Runge-Kutta integrators. We leave the development of higher-order methods for future work.
The simulation runs with the adiabatic EOS (for both the Roe and the HLLC solvers) exhibit systematically higher ($\sim$ an order of magnitude) error levels, with first-order convergence. This pattern is likely attributed to the complexities the adiabatic EOS introduces to the system, including the additional energy equation to be solved. Additionally, we note that our simulations may encounter stability issues in scenarios involving supersonic adiabatic flows with added perturbations (white noise). Future work will address these challenges and improve the handling of adiabatic processes within our simulations.
Figure 9: Time-averaged L2 norm relative error for diagonal sound wave test, as a function of simulation resolution (cells per unit length). Overall, convergence between first- and second-order is achieved (see the dashed lines for reference).

### v2

@@ 4.2.1 Horizontal sound waves with uniform collapse
We first considered a purely horizontally propagated traveling wave with a uniform collapse profile. This can also provide a comprehensive view of the wave behavior in the local box. We set up the following perturbation:
@@
where $A_{0}$ is some (real) constant, and $c_{\mathrm{s},\mathrm{H0}}=c_{\mathrm{s0}}/R_{0}$ is the initial value of the effective horizontal sound speed defined in Sect. 2. This perturbation corresponds to wave amplitudes of $\left(X,\Pi\right)=\left(\frac{iA_{0}}{2c_{\mathrm{s}}},\frac{A_{0}k_{x}}{2R_{0}}\right)$.
For the numerical test of the local model, we chose a collapse profile with $R_{0}=10.0L_{0}$, $U_{0}=1.0c_{\mathrm{s}0}$, $L_{z0}=1.0L_{0}$, and $U_{R0}=0$, with a perturbation amplitude of $A_{0}=10^{-6}$, $k_{x}=16\pi$. The simulation box size was set as $L_{x,\mathrm{sim}}=0.5L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 256 cells per unit length.
The relative amplitude of gas density perturbation, $\delta\rho/\rho$, was measured with respect to the background gas density, $\rho$, at each time point, which was obtained by averaging the gas density over the domain. The space-time plot of the vertically averaged value of $\delta\rho/\rho$ presented in the upper panel of Fig. 5 shows the evolution of the wave pattern. The wave propagates rightward with a time-varying effective sound speed, $c_{\mathrm{s},\mathrm{H}}=c_{\mathrm{s}0}/\mathcal{R}$. As the collapse proceeds, the wave propagation is accelerated as $\mathcal{R}$ decreases, matching the overlaid curves of theoretical phase shift profiles over time. This time-varying effective sound speed is not physical in the global picture, but is instead due to our choice of a time-dependent reference frame in the local model.
Our choice of collapse profile results in the wave operating in the WKB regime. It is thus appropriate to compare the evolution of the wave amplitude to that expected from the WKB approximation solution. The lower panel of Fig. 5 shows this comparison of the relative amplitude of density perturbation, $A_{\rho}=\max\left(\delta\rho/\rho\right)$. The numerical result of the $A_{\rho}$ evolution agrees well with the WKB approximation solution, $A_{\rho}=A_{\rho,0}\mathcal{R}^{-1/2}$.
Similarly, for a purely vertical sound wave, the corresponding length scale becomes $L_{z}$ instead of $\mathcal{R}$. We have also numerically tested the vertical sound wave and found that the effective vertical sound speed evolves following $c_{s,H}=c_{s0}/{L_{z}}$, and the relative amplitude of $\delta\rho$ follows $A_{\rho}=A_{\rho,0}L_{z}^{-1/2}$, which aligns with the theoretical expectations.
4.2.2 Horizontal waves with nonuniform collapse
We next tested the horizontal sound waves with a more general power law collapse profile. In order to compare with the exact solution Eq. (90-92), we considered a collapse profile with $\beta=1/2$ and $t_{\mathrm{c}}=1$, i.e., $\mathcal{R}={R}_{0}(1-t)^{1/2}$, and correspondingly initiated the perturbation of velocity and density as
@@ 4.2.2 Horizontal waves with nonuniform collapse
where according to eq. (93), $s_{0}=s(0)=\left(t_{c}/t_{0}\right)^{1-\beta}$, and $t_{0}$ can again be obtained from eq. (93). $A_{0}$ is some real constant, and this perturbation corresponds to $A=0$, $B=A_{0}/2$ in eq. (90-92). For the numerical test, we set $R_{0}=1.0L_{0}$, $L_{z0}=1.0L_{0}$, $U_{R0}=0$, with a perturbation of $A_{0}=10^{-5}$, $k_{x}=2\pi$, and a simulation box size of $L_{x,\mathrm{sim}}=2.0L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 256 cells per unit length.
Similar to Sect. 4.2.1, the relative amplitude of the gas density perturbation, $\delta\rho/\rho$, was measured at each time point, where $\rho$ was obtained by averaging the gas density over the domain. The space-time plot of the vertically averaged relative gas density, $\delta\rho/\rho$, displayed in the upper panel of Fig. 6 shows that this perturbation leads to a horizontal standing wave. This wave exhibits phase reversal with each half wave period. We further measured the relative amplitude of the density perturbation by performing a Fourier transform to the wave pattern, and identified the amplitude of the dominant frequency component. The time evolution of the relative amplitude, $\delta\rho/\rho$, is shown in the lower panel of Fig. 6, with the length scale corresponding to the freeze-out regime, $L_{\mathrm{freezeout}}$ (see Eq. (94)), marked with a vertical dotted line. The result is in good agreement with the analytical solution in both the WKB regime ($\mathcal{R}>L_{\mathrm{freezeout}}$) and the freeze-out regime ($\mathcal{R}<L_{\mathrm{freezeout}}$). The L2 norm of the relative error for the relative amplitude, $\mathrm{err}(\delta\rho/\rho)$, defined in eq. (70) is on the order of $10^{-5}$ when the absolute value of the amplitude solution is much larger than 00. When the solution of the amplitude is close to 00 (i.e., close to the period of phase reversal), the relative error, $\mathrm{err}(\delta\rho/\rho)$, can be up to the order of $10^{-2}$, partly due to the division of a small analytical solution value, and partly due to inherent limitations and numerical inaccuracies that can arise near the phase reversal period. The slightly diffused wave pattern during the phase reversal can also be seen in Fig. 6.
Figure 7: Test case for diagonal sound wave. Upper panel: Illustration of the diagonal sound wave and the shear flow generated. Lower panel: Amplitude of the shear flow generated by the wave as a function of R. The numerical result (solid red line) closely matches the analytical solution (dashed black line).
@@ 4.2.3 Diagonal sound waves
Diagonally propagating sound waves may introduce numerical complexities, as they generate shear flows perpendicular to the direction of wave propagation. This is illustrated in the upper panel of Fig. 4, with the wave pattern characterized by the density.
We tested diagonal sound waves by generating density and velocity perturbations of
@@
where $A_{0}$ is some (real) constant. This perturbation corresponds to an initial wave amplitude of $\left(X_{0},\Pi_{0}\right)=\left(\frac{iA_{0}}{2c_{\mathrm{s}0}},\frac{A_{0}}{2}\sqrt{k_{x}^{2}R_{0}^{-2}+k_{z}^{2}L_{z0}^{-2}}\right)$. Recalling Eq. (80), Eq. (81), and Eq. (95), this pure diagonal sound wave generates a shear flow that can be characterized by
@@
and the dynamics of the system of $(X,\Pi)$ can be obtained from Eqs. (80) and (81). Thus, the amplitude $X$ can be obtained by solving the equation
@@
where the wave frequency, $\omega$, is given in Eq. (82).
We numerically tested the diagonal sound wave by setting $R_{0}=1.0L_{0}$, $U_{0}=1.0c_{\mathrm{s}0}$, $L_{z0}=1.0L_{0}$, and $U_{R0}=0$, with a perturbation of $A_{0}=10^{-5}$, $k_{x}=2\pi$, $k_{z}=4\pi$, and a simulation box of $L_{x,\mathrm{sim}}=2.0L_{0}$, $L_{z,\mathrm{sim}}=1.0L_{0}$, with a resolution of 512 cells per unit length.
In the lower panel of Fig. 7, we compare the amplitude of the shear flow generated by the diagonal sound wave in the numerical simulation with the analytical solution. The numerical result is obtained by calculating the shear flow, $\delta v_{\mathrm{shear}}$, in each grid, and performing a 2D Fourier transform of the shear flow map, followed by identifying the amplitude of the dominant frequency component. The analytical solution of the shear flow amplitude, $2{c_{\mathrm{s}}\mathcal{R}^{-2}}k_{x}k_{z}\left(1-b^{-2}\right)X$, was obtained by numerically solving for the wave amplitude, $X$, at each time step according to Eq. (108), with the explicit Runge-Kutta integration method of order 5(4) (Dormand80). The numerical result and analytical solution of the shear flow amplitude are in good agreement. We further measured the L2 norm of the relative error for the shear strength, $\delta v_{\mathrm{shear}}$, in each grid, and got $\mathrm{err}(\delta v_{\mathrm{shear}})$ as defined in Eq. (70) of the order of $10^{-6}$.
Figure 8: Conservation laws for tests of horizontal shear flow (left) and of white noise in the velocity field (right). For each case, the normalized variation in $J\bar{\rho}$ (left), $\Pi_{i}$ (middle), and $H_{k}$ (right) is presented as a function of time. For the horizontal shear flow test (see also Sect. 4.1.1), the modified mass conservation law ($J\bar{\rho}$) and the conservation law for total momenta ($\Pi_{i}$) and for kinetic helicity $H_{k}$ are all satisfied to machine precision (note that for all of the panels, the y axis ticks are in units of $10^{-13}$). For the case with white noise in the velocity field, the conservation laws for both mass, $J\bar{\rho}$, and total momenta, $\Pi_{i}$, are fulfilled to machine precision as well, but not for the kinetic helicity, $H_{k}$ (note that for the upper two panels, the y axis ticks are in units of $10^{-14}$). For comparison, a simulation with the same setup of white noise, but without collapse (i.e., $U_{0}=U_{R0}=0$) was performed (“WN0”, dashed grey lines, see also the legend), which gives consistent results to WN. We thus anticipate that the deviation from conservation law for $H_{k}$ is likely due to intrinsic characteristics of Athena++, which in certain situations is not strictly vorticity-conserving.
@@ 4.3 Conservation laws
The continuity equation of the local model, Eq. (5), gives the conservation law for the total mass,
@@
The momentum equation of the local model, Eq. (6 - 8), gives the conservation law for the total momenta,
@@
Another conserved quantity in the local model is the kinetic helicity, $H_{k}$, of the relative flow, which is a quantity associated with the vorticity. It is defined as
@@
In order to validate these conservation laws in the local box, we performed a simulation run with a uniform background density, $\rho_{0}=1$, introducing white noise perturbations in the velocity field, $v_{x},v_{y}$, and $v_{z}$, with a relative amplitude of $10^{-6}$. We chose a collapse profile with $R_{0}=10.0$, $U_{0}=-1.0c_{\mathrm{s}0}$, $L_{z0}=10.0$, and $U_{R0}=-0.1$, and set the simulation box size as $L_{x,\mathrm{sim}}=L_{y,\mathrm{sim}}=0.5$, $L_{z,\mathrm{sim}}=1.0$, with a resolution of 32 cells per unit length. This test run is labeled as “WN.” For comparison, we also performed a simulation with the same setup for the white noise, but without collapse (i.e., $U_{0}=U_{R0}=0$), which is labeled as “WN0.” In addition, the result of the horizontal shear flow test (see Sect. 4.1.1) was also used to examine the conservation laws.
We measured the volume-averaged density, $\bar{\rho}$ , total momenta, $\Pi$, and kinetic helicity, $H_{\rm k}$, in the simulations, and present in Fig. 8 the normalized variation in $J\bar{\rho}$ , $\Pi$, and $H_{\rm k}$ as a function of time. Our results demonstrate that the modified mass conservation law ($J\bar{\rho}$) and the conservation law for total momenta ($\Pi_{i}$) are both satisfied to machine precision. The kinetic helicity conservation, although preserved to machine precision for the shear flow test, exhibits significant deviation for the white noise simulations. This discrepancy persists in both WN and WN0 cases, suggesting that the lack of conservation is not a consequence of the integration of the collapsing box. Instead, it is likely due to intrinsic characteristics of the Athena++ framework, which in certain situations is not strictly vorticity conserving, likely because the higher-order Godunov scheme generally introduces a degree of numerical diffusion (see e.g. Seligman17). In the white noise case in which vorticity is dominated by fine-scale structures, there is increased numerical dissipation at small scales because the slope limiter is expected to lead to the nonconservation of $H_{\rm k}$. For an adiabatic EoS, the entropy conservation law is also validated; this is further presented in App. C.
@@ 4.4 Convergence study
We performed a convergence study with the diagonal wave (Sect. 4.2.3) test, as it shows the maximum possible error. Basic simulation setups such as wave properties and box sizes are the same as presented before. Fiducially, simulations were performed with Roe’s solver, the VL2 integrator, PPM reconstruction, and an isothermal EoS. The CFL number in the convergence study was fixed to $0.2$. However, we validated the numerical stability of our code for CFL numbers up to 0.4 for VL2 and 0.8 for RK3. We further explored the effects of different Riemann solvers, time integrators, spatial reconstruction methods, and the EoS on the numerical performance of the code. We also varied the spatial resolution of our simulation in units of cells per unit length. The simulation setups for test runs that we performed for the convergence study are presented in Table 1. Figure 9 shows the time-averaged L2 relative error, $\mathrm{err}(\delta v_{\mathrm{shear}})$, as a function of resolution.
Table 1: Simulation setup for the convergence study.
@@
 [^3] Variations compared to the fiducial simulation highlighted in bold. “Reconst.” stands for reconstruction.
Our results demonstrate that our code generally achieves convergence between the first and second orders for the diagonal wave test. The performances for test cases HLLE and RK3 remain generally similar to the fiducial case, while a lower order of spatial reconstruction (PLM) leads to a lower accuracy at lower resolutions. These tests with an isothermal EoS achieve second-order convergence at lower resolutions, with a tentative trend toward the first order at high resolutions. This trend is likely attributed to time-dominated errors at higher resolutions, as our source term incorporation is first-order in time. We also explored the higher-order method to integrate source terms with the VL2 integrator, which was not as efficient as our first-order method, showing a comparable or slightly higher error, and reduced stability for adiabatic flows. Therefore, we chose to use our first-order method. The source term can also be incorporated into higher-order methods with Runge-Kutta integrators. We leave the development of higher-order methods to future work.
The simulation runs with the adiabatic EoS (for both the Roe and the HLLC solvers) exhibit systematically higher ($\sim$ an order of magnitude) error levels, with first-order convergence. This pattern is likely attributed to the complexities that the adiabatic EoS introduces to the system, including the additional energy equation to be solved. Additionally, we note that our simulations may encounter stability issues in scenarios involving supersonic adiabatic flows with added perturbations (white noise). Future work will address these challenges and improve the handling of adiabatic processes within our simulations.
Figure 9: Time-averaged L2 norm relative error for diagonal sound wave test, as a function of simulation resolution (cells per unit length). Overall, convergence between the first and second orders is achieved (see the dashed lines for reference).